OpenStudy (anonymous):

CORRECT ME Find the volume of a sphere with surface area of 452.39cm2. Round your answer to the nearest hundredth. A. 150.79 cm3 B. 904.8 cm3 C. 904.78 cm3 D. 7238.2 cm3

OpenStudy (anonymous):

V=(r/3)surface area r=√(surface area/4π) r=9 cm V=(452.39)(9)/3 V=904.78 cm³

hartnn (hartnn):

i get r = 6 try again ?

OpenStudy (primeralph):

r = 6

hartnn (hartnn):

r=√(surface area/4π) = r=√(452.39/4π) =...?

OpenStudy (primeralph):